OpenStudy (anonymous):

Solve for y, then find the value of y when given x = –2 6x = 7 - 4y

OpenStudy (jlastino):

You should subtract subtract 7 from both sides of the equation first to equate it yo y then you can solve for y by replacing x with -2

OpenStudy (anonymous):

i can figure it out : p lol

OpenStudy (jlastino):

You subtract 7 from both sides 6x-7 = 7-4y-7 It would become like this 6x-7 = -4y Then replace x with -2 6(-2)-7=-4y Can you solve for y from here?

OpenStudy (anonymous):

no : o

OpenStudy (jlastino):

-12-7=-4y -19=-4y Dvide -19 by -4. what do you get?

OpenStudy (anonymous):

like 4 and 1/3